Masongray25:

if a bedroom is rectangular and the perimeter is 44 feet and the length is twice teh width what is the dimensions

supie:

"The length twice the width" meaning that you have to find the length then when you have the length multiply it by 2 then you'll have the width so \(\mathbb L*2=W\)

supie:

@Masongray25

jhonyy9:

@supie i think invers l = 2W

jhonyy9:

perimeter = 44 P = 2(l +w) P = 2(2w +w) P = 2*3w P = 6w 44 = 6w w = 44/6 = 7,3 l = 2w = 2*7,3 = 14,6

jhonyy9:

@Laylalyssa